# An Incremental Unified Framework for Small Defect Inspection

### v2

@@ Abstract
Artificial Intelligence (AI)-driven defect inspection is pivotal in industrial manufacturing. Yet, many methods, tailored to specific pipelines, grapple with diverse product portfolios and evolving processes. Addressing this, we present the Incremental Unified Framework (IUF), which can reduce the feature conflict problem when continuously integrating new objects in the pipeline, making it advantageous in object-incremental learning scenarios. Employing a state-of-the-art transformer, we introduce Object-Aware Self-Attention (OASA) to delineate distinct semantic boundaries. Semantic Compression Loss (SCL) is integrated to optimize non-primary semantic space, enhancing network adaptability for novel objects. Additionally, we prioritize retaining the features of established objects during weight updates. Demonstrating prowess in both image and pixel-level defect inspection, our approach achieves state-of-the-art performance, proving indispensable for dynamic and scalable industrial inspections.
Our code will be released at https://github.com/jqtangust/IUF.
@@
Small Defect Inspection; Incremental Unified Framework; Incremental Learning
@@ 1 Introduction
AI-driven small defect inspection, commonly known as anomaly detection, is pivotal in numerous industrial manufacturing sectors [12, 22, 38], spanning from medical engineering [15] to material science [14], and electronic components [13]. Automated product inspections facilitated by these applications not only boast impressive accuracy but also significantly reduce labor costs.
@@
Figure 2: Semantic space in our methodology (B-D). In (A), all objects in the original semantic space are tightly coupled, leading to “catastrophic forgetting", i.e., the learning of new objects will result in forgetting previously learned objects. Our method firstly builds a semantic boundary of each object (B), then compacts the non-primary semantic space (C), and finally suppresses semantic updates of the primary object’s features (D).
Currently, only limited research addresses these challenges. One-Model-N-Object approaches like UniAD [34] and OmiAD [37] propose a unified network suitable for a broad spectrum of objects. However, its capability is restricted when encountering dynamic or unfamiliar objects. On the other hand, CAD [18] suggests embedding features of multiple objects into a memory bank, with a binary classifier discerning defects (Fig. 1 (D)). Yet, the differences in object distributions can jeopardize memory bank features, escalating storage demands and impeding pixel-level performance evaluation.
In an attempt to surmount the confines of prior research, we introduce an Incremental Unified Framework (IUF), which integrates the multi-object unified model with object-incremental learning, as depicted in Fig. 1 (E). IUF capitalizes on the unified model’s prowess, enabling pixel-precise defect inspection across diverse objects without necessitating an embedded feature storage in a memory bank.
In this framework, one significant challenge is “catastrophic forgetting", which hampers network efficacy in the Incremental Unified framework due to semantic feature conflicts in the reconstruction network (Fig. 2 (A)). We propose Object-Aware Self-Attention (OASA), leveraging object category features to segregate the semantic spaces of different objects (Fig.2 (B)). These features serve as semantic constraints within the encoding process of the reconstruction network, establishing distinct feature boundaries in the semantic hyperplane, thereby alleviating feature coupling. Then, we introduce Semantic Compression Loss (SCL) to condense non-essential feature values, concentrating the feature space on principal components (Fig.2 (C)). This creates more semantic space for optimizing unseen objects and minimizing potential forgetting issues. Finally, when learning novel objects, we propose a new updating strategy to retain features of prior objects and reduce interference from new objects in the prevailing feature space (Fig.2 (D)). Our framework offers a versatile solution for small defect inspection, adeptly navigating the challenges posed by feature space conflicts, and achieves state-of-the-art (SOTA) performance at both image and pixel levels.
@@
  • We propose the Incremental Unified Framework (IUF) that effectively deals with small defect inspection in real industrial scenarios, which is particularly suitable for practical systems with dynamic and frequently-adjusted products.
  • With the Object-Aware Self-Attention (OASA), Semantic Compression Loss (SCL), and the new updating strategy, our framework reduces the potential risk of feature conflicts and mitigates the catastrophic forgetting issue.
  • With our Incremental Unified framework, our method can provide not only image-level performance but also pixel-level location. Compared to other baselines, our method achieves SOTA performance.
@@ 2 Related Work
2.1 Small Defect Inspection
Small defect inspection primarily falls into two paradigms: feature embedding-based and reconstruction-based methods. Both aim to pinpoint the locations of defects.
Feature embedding-based methods localize defective regions by discerning feature distribution discrepancies between defective and non-defective images, e.g., SPADE [9], PaDim [10], PatchCore [26], GraphCore [32], SimpleNet [20]. These approaches extract diverse image features using a deep network and subsequently assess the dissimilarities between test image features and those of the training dataset. If these disparities are substantial, they indicate defective regions.
Another more efficient method is the reconstruction-based approach that operates on an alternative hypothesis. Such methods posit that models trained on normal samples excel in reproducing normal regions while struggling with anomalous regions [4, 8, 34, 37]. Therefore, many approaches aim to discover an optimal network architecture for effective sample reconstruction, including Autoencoder (AE) [4], Variational Autoencoder (VAE) [19], Generative Adversarial Networks (GANs) [1], Transformer [34], etc. Furthermore, beyond image-level reconstruction, certain methods endeavor to reconstruct features at a more granular level, e.g., InTra [24] and UniAD [34]. Until recently, the majority of prior approaches adhered to the One-Model-One-Object pattern, as shown in Fig. 1 (A). However, recent advancements such as UniAD [34] and OmniAL [37] have extended the reconstruction paradigm to the One-Model-N-Objects pattern (depicted in Fig. 1 (C)). This progress curtails memory usage compared to the one-model-one-object approach, propelling the entire problem into a more general framework.
@@ 2.1 Small Defect Inspection
2.2 Incremental Learning for Defect Inspection
Incremental learning (also known as continuous learning, lifelong learning) aims to enable AI systems to continuously acquire, update, accumulate, and utilize knowledge in response to external changes [30]. There have been some previous studies combining incremental learning and small defect inspection, but these studies mainly focused on augmenting the defect types of one object under one specific scenario [33, 5, 28, 27, 7]. The purpose of these methods is to improve the performance of the model. Most recently, CAD [18] initially proposed an unsupervised feature embedding-based method for continuous learning of task sequences, but there are several problems.
First, due to the existence of significant differences in the distribution of objects, the algorithm uses a single memory bank to store all the object information, which will have interference between objects. In addition, as continuous learning proceeds, the memory bank gradually expands, which puts a lot of pressure on the storage space. Finally, since the reference feature space cannot be exactly the same as the detected objects, this algorithm cannot compute the pixel-level performance, but can only give the image-level performance.
Based on this, we restate the task and integrate an incremental unified framework to overcome the above problems. We propose the first reconstruction-based model for this continuous learning scenario. This method can effectively avoid the problem of feature space distribution differences between different objects, and only needs a reasonable feature space reassignment in reconstruction.
@@ 4 Methodology
To mitigate catastrophic forgetting, we propose three critical components for Incremental Unified small defect inspection. Firstly, we identify the semantic boundary to highlight the differences among categories, thereby facilitating the preservation of old knowledge (Sec. 4.1). Secondly, we compact the semantic space to maximize the compression of redundant features, which helps to alleviate the issue of semantic conflicts (Sec. 4.2). Finally, we sustain primary semantic memory to minimize the feature forgetting of the old object, thereby improving the retention of critical knowledge (Sec. 4.3).
@@ 4.1 Identifying Semantics Boundary
Previous unified approaches [34] directly reconstruct all objects through one autoencoder, as in Fig. 4 (A), so multiple objects share the same semantic space in one model. As a result, the semantic spaces of multiple objects are tightly coupled.
However, the tight semantic coupling is likely to cause undifferentiated updating of the feature space of the old objects when learning a new object, consequently resulting in a significant catastrophic forgetting of the old knowledge. Therefore, we introduce a more explicit constraint to facilitate the network to distinguish the feature semantic space among different objects. Specifically, for step $n$, we introduce its classifying label, $L_{n}$, as in Eq. (3),
@@
where $x^{+}_{o_{n}}\in\mathbb{R}^{C\times H\times W}$ is the normal images in current step, $y_{n}$ is the final output of the discriminator, $D(\cdot)$, and $\sigma$ is the network parameter in this discriminator. $L_{CE}(\cdot)$ is the cross-entropy loss, which contains $N$ objects. Besides, $D(\cdot)$ also outputs $C_{o_{n}}\in\mathbb{R}^{T\times C\times H\times W}$, which corresponding to $T$ key layers of semantic features, as Fig. 4 (B). By introducing the classifying label, the semantic space of each object is restricted to its corresponding classification, which helps to identify the semantic boundary between different objects.
@@
$\text{Attention}(C_{o_{n}},Q,K,V)=\texttt{softmax}(\frac{(C_{o_{n}}\cdot Q)K^{T}}{\sqrt{d_{k}}})V,$ (4)
where $\cdot$ is Hadamard product, $d_{k}$ is the dimension of the key or query, and $T$ is the matrix transpose operator. $Q\in\mathbb{R}^{T\times C\times H\times W}$ is the query in $T$ key layers of the transformer-based network. $K$ and $V$ are the keys and values of this network, respectively. By inserting $C_{o_{n}}$ to $Q$, the reconstruction network can identify semantic boundaries explicitly.
@@ 4.2 Compacting Semantic Space
By identifying semantic boundaries, we can further compress the feature space of old objects as much as possible, thus reserving more network capacity for new objects, which can reduce feature conflicts. To meet this goal, we design a compact feature hyperplane regularization, which helps us to eliminate the redundant features of old objects and leave more capacity for future unknown new objects.
Specifically, the latent features $M\in\mathbb{R}^{B\times C\times H\times W}$ is firstly compressed by aggregating spatial features as in Eq. (5),
@@
$S_{n}$ reflects the importance of semantic information for $M$ on each channel dimension. Larger eigenvalues correspond to more important semantic information, while smaller eigenvalues correspond to relatively non-primary semantic information. Based on this fact, we can compress the features of each object by reducing the non-primary semantic information. Therefore, we construct the Semantic Compression Loss (SCL) as in Eq. (7) and Fig. 5,
@@
where $t\in[1,C]$ is a hyperparameter, which is a scale factor that reflects the degree of compression of the semantic feature space. To construct the total loss, we use $L=\lambda_{0}L_{1}+\lambda_{1}L_{CE}+\lambda_{2}L_{sc}$, where $\lambda_{0}=1$, $\lambda_{1}=0.5$ and $\lambda_{2}$ is hyperparameter for total loss. $\lambda_{2}$ is set between $1-10$ based on the experiment. This loss continuously compacts the non-primary feature space while achieving the final optimization goal.
@@ 5 Experiments
In this section, we first discuss the experimental setup, including the dataset, baselines, and protocol. Then, we conduct experiments under these protocols, showing quantitative and qualitative experimental results. Finally, we conduct ablation studies to analyze the impact of our proposed strategy on the overall performance.
@@
Table 1: Quantitative evaluation in MvTec AD [3] (A) and VisA [39] (B). “Image / Pixel" shows the image-level and pixel-level performance respectively. “NA" is Not Available, since CAD [18]-based method cannot locate defects in pixel-level. Red and Gray Red represents the best image-level and pixel-level performance respectively.
5.1 Experimental Setup
[…]
@@ 5.1 Experimental Setup
Figure 6: Qualitative Evaluation. Compared with other baselines, our method can significantly reduce semantic feature conflict and locate defects more accurately.
@@ 5.3 Ablation Study
Table 2: Quantitative evaluation in ablation study. The results show that all components in our method, Object-Aware Self-Attention, semantic compression loss, and our updating strategy, contribute to performance improvement of the whole framework. Red and Gray Red represents the best image-level and pixel-level performance, respectively.
Effectiveness of Object-Aware Self-Attention To demonstrate the effectiveness of classifying attributes, we ablate Object-Aware Self-Attention in our structure. “w/o OASA” in Table 2 shows that this structure can reduce catastrophic forgetting by classifying information insertion. Moreover, in Fig. 7 (“w/o OASA”), if there is no semantic boundary, it will lead to serious feature conflict, and then impact the performance of the reconstruction network.
Effectiveness of Semantic Compression Loss To demonstrate the effectiveness of Semantic Compression Loss, we ablate this loss in our training. “w/o SCL” in Table 2 shows that this structure can reduce catastrophic forgetting by preserving more space for new objects. Besides, in Fig. 7 (“w/o SCL”), by compressing the space, we obtain clearer defect localization, which indicates that more network space can be obtained for incremental learning.
Effectiveness of Our Updating Strategy To demonstrate the effectiveness of our updating strategy, we ablate this updating method, and “w/o US” in Table 2 shows that this method can reduce catastrophic forgetting. Also, in Fig. 7 (“w/o US”), our algorithm can help memorize important semantic spaces for old objects, thus reducing the rewriting of semantic space in the reconstruction network.
@@
Figure 7: Qualitative evaluation of ablation study. Compared with “w/o US” and “w/o OSOA”, it is obvious that our method can be more accurate to locate defects. Compared with “w/o SCL”, since increasing the available space of the model, our algorithm can further reduce the feature conflict and reduce more interference in the background (Please zoom in for more details).
6 Limitations
Firstly, our approach does not explicitly represent the semantic space capacity, which is unfavorable for measuring the network ability for object-incremental learning. Besides, feature sharing is also significant in incremental learning since feature sharing on the reconstruction network can reduce the space occupation of individual objects. These limitations are valuable and should be further discussed.

### v3

@@ Abstract
Artificial Intelligence (AI)-driven defect inspection is pivotal in industrial manufacturing. However, existing inspection systems are typically designed for specific industrial products and struggle with diverse product portfolios and evolving processes. Although some previous studies attempt to address object dynamics by storing embeddings in the reserved memory bank, these methods suffer from memory capacity limitations and object distribution conflicts. To tackle these issues, we propose the Incremental Unified Framework (IUF), which integrates incremental learning into a unified reconstruction-based detection method, thus eliminating the need for feature storage in the memory. Based on IUF, we introduce Object-Aware Self-Attention (OASA) to delineate distinct semantic boundaries. We also integrate Semantic Compression Loss (SCL) to optimize non-primary semantic space, enhancing network adaptability for new objects. Additionally, we prioritize retaining the features of established objects during weight updates. Demonstrating prowess in both image and pixel-level defect inspection, our approach achieves state-of-the-art performance, supporting dynamic and scalable industrial inspections.
Our code is released at https://github.com/jqtangust/IUF.
@@
Small Defect Inspection; Incremental Unified Framework
@@ 1 Introduction
AI-driven small defect inspection, commonly known as anomaly detection, is pivotal in numerous industrial manufacturing sectors [14, 26, 42, 36], spanning from medical engineering [17] to material science [16], and electronic components [15]. Automated product inspections facilitated by these applications not only boast impressive accuracy but also significantly reduce labor costs.
@@
Figure 2: Semantic space in our methodology (B-D). In (A), all objects in the original semantic space are tightly coupled, leading to “catastrophic forgetting”, i.e., the learning of new objects will result in forgetting previously learned objects. Our method firstly builds a semantic boundary of each object (B), then compacts the non-primary semantic space (C), and finally suppresses semantic updating in the previous objects’ feature space (D).
Currently, only limited research addresses these challenges. One-Model-N-Object approaches like UniAD [38] and OmiAD [41] propose a unified network suitable for a broad spectrum of objects (Fig. 1 (C)). However, its capability is restricted when encountering dynamic or unfamiliar objects. On the other hand, CAD [20] suggests embedding features of multiple objects into a memory bank, with a binary classifier discerning defects (Fig. 1 (D)). Yet, the differences in object distributions can jeopardize memory bank features and impede pixel-level performance evaluation. Although Liu et al. [22] design a learnable prompt to find needed object information in the memory bank, this solution is still constrained by escalating storage demands.
To surmount the confines of prior research, we introduce an Incremental Unified Framework (IUF), which integrates the multi-object unified model with object-incremental learning, as depicted in Fig. 1 (E). IUF capitalizes on the unified model’s prowess, enabling pixel-precise defect inspection across diverse objects without necessitating an embedded feature storage in the memory bank.
In this framework, one significant challenge is “catastrophic forgetting”, which hampers network efficacy in the Incremental Unified framework due to semantic feature conflicts in the reconstruction network (Fig. 2 (A)). We propose Object-Aware Self-Attention (OASA), leveraging object category features to segregate the semantic spaces of different objects (Fig.2 (B)). These features serve as semantic constraints within the encoding process of the reconstruction network, establishing distinct feature boundaries in the semantic hyperplane, thereby alleviating feature coupling. Then, we introduce Semantic Compression Loss (SCL) to condense non-essential feature values, concentrating the feature space on principal components (Fig.2 (C)). This creates more semantic space for optimizing unseen objects and minimizing potential forgetting issues. Finally, when learning novel objects, we propose a new updating strategy to retain features of prior objects and reduce interference from new objects in the prevailing feature space (Fig.2 (D)). Our framework offers a versatile solution for small defect inspection, adeptly navigating the challenges posed by feature space conflicts, and achieves state-of-the-art (SOTA) performance at both image and pixel levels.
@@
  • We propose the Incremental Unified Framework (IUF), which can dynamically detect small defects in frequently-adjusted industrial scenarios. To the best of our knowledge, IUF is the first framework to integrate incremental learning into the unified reconstruction-based small defect detection method, thereby overcoming memory bank capacity limitations and object distribution conflicts.
  • Our framework incorporates Object-Aware Self-Attention (OASA), Semantic Compression Loss (SCL), and a new updating strategy, which together reduce the potential risk of feature conflicts and mitigate the catastrophic forgetting issue.
  • The Incremental Unified Framework (IUF) enables our method to deliver not only image-level performance but also pixel-level location. Compared to other baselines, our method achieves SOTA performance.
@@ 2 Related Work
Small Defect Inspection Small defect inspection (Anomaly Detection) primarily falls into two paradigms: feature embedding-based and reconstruction-based methods. Both aim to pinpoint the locations of defects.
Feature embedding-based methods localize defective regions by discerning feature distribution discrepancies between defective and non-defective images, e.g., SPADE [9], PaDim [10], PatchCore [30], GraphCore [35], SimpleNet [24]. These approaches extract diverse image features using a deep network and subsequently assess the dissimilarities between test image features and those of the training dataset. If these disparities are substantial, they indicate defective regions. Another efficient method is the reconstruction-based approach, which relies on an alternative hypothesis. Such methods posit that models trained on normal samples excel in reproducing normal regions while struggling with anomalous regions [4, 8, 38, 41]. Therefore, many approaches aim to discover an optimal network architecture for effective sample reconstruction, including Autoencoder (AE) [4], Variational Autoencoder (VAE) [23], Generative Adversarial Networks (GANs) [1], Transformer [38], etc. Furthermore, beyond image-level reconstruction, certain methods endeavor to reconstruct features at a more granular level, e.g., InTra [28] and UniAD [38]. Until recently, the majority of prior approaches adhered to the One-Model-One-Object pattern, as shown in Fig. 1 (A). However, recent advancements such as UniAD [38] and OmniAL [41] have extended the reconstruction paradigm to the One-Model-N-Objects pattern (depicted in Fig. 1 (C)). This progress curtails memory usage compared to the one-model-one-object approach, propelling the entire problem into a more general framework.
@@
Incremental Learning for Small Defect Inspection There have been some previous studies combining incremental learning and small defect inspection; however, these studies mainly focused on augmenting the defect types of one object under a specific scenario [37, 5, 32, 31, 7] (Fig. 1 (C)). The primary goal of these methods is to enhance the model’s performance, which does not align with the requirements of object-incremental learning.
Recently, Li et al.[20] initially proposed a feature embedding-based method for continuous learning of task sequences. However, due to the algorithm using a single memory bank to store all object information, which leads to interference between objects, the reference feature space cannot be exactly the same as the detected objects. Consequently, this algorithm cannot compute pixel-level performance and can only provide image-level performance. Most recently, Liu et al.[22] proposed an unsupervised continual anomaly detection (UCAD) method. Although it attempts to distinguish features of different objects using contrastive learning, this strategy still requires the assistance of a memory feature bank.
Unlike previous solutions, our method avoids interference from multiple object classes caused by explicit memory objects. We propose the first reconstruction-based model for this continuous learning problem. Therefore, our method can effectively avoid the problem of feature space distribution differences between different objects and only requires a reasonable feature space reassignment in the reconstruction network.
[…]
@@ 4 Methodology
Overview To mitigate catastrophic forgetting, we propose three critical components for Incremental Unified small defect inspection. Firstly, we identify the semantic boundary to highlight the differences among categories, thereby facilitating the preservation of old knowledge (Sec. 4.1). Secondly, we compact the semantic space to maximize the compression of redundant features, which helps to alleviate the issue of semantic conflicts (Sec. 4.2). Finally, we sustain primary semantic memory to minimize the feature forgetting of the old object, thereby improving the retention of critical knowledge (Sec. 4.3).
@@ 4.1 Identifying Semantics Boundary
Previous unified method [38] directly reconstructs all objects through one autoencoder, as in Fig. 4 (A), so multiple objects share the same semantic space in one model. So, the semantic spaces of different objects are tightly coupled.
This tight semantic coupling is likely to cause undifferentiated updating of the feature space of the old objects when learning a new object, consequently resulting in a significant catastrophic forgetting of the old knowledge. Therefore, we introduce a more explicit constraint to facilitate the network to distinguish the feature semantic space among different objects. Specifically, for step $n$, we introduce its category label, $L_{n}$, as in Eq. (3),
@@
where $x^{+}_{o_{n}}\in\mathbb{R}^{3\times H_{o}\times W_{o}}$ indicates the normal object images in current step, $y_{n}$ is the final output of the discriminator, $D(\cdot)$, and $\sigma$ is the parameter in this discriminator. $L_{CE}(\cdot)$ is the cross-entropy loss, which contains $N$ objects. Besides, $D(\cdot)$ also outputs $C_{o_{n}}\in\mathbb{R}^{T\times C\times H\times W}$, which corresponding to $T$ key layers of semantic features, as Fig. 4 (B). By introducing the category label, each object is limited to its corresponding semantic space in the reconstruction network, which helps to identify the semantic boundary between different objects.
@@
$\text{Attention}(C_{o_{n}},Q,K,V)=\textnormal{softmax}(\frac{(C_{o_{n}}\cdot Q)K^{T}}{\sqrt{d_{k}}})V,$ (4)
where $\cdot$ is Hadamard product, $d_{k}$ is the dimension of the key or query, and $T$ is the matrix transpose operator. $Q\in\mathbb{R}^{T\times C\times H\times W}$ is the query in $T$ key layers of the transformer-based network. $K$ and $V$ are the keys and values of this network, respectively. By inserting $C_{o_{n}}$ to $Q$, the reconstruction network can explicitly identify semantic boundaries.
@@ 4.2 Compacting Semantic Space
By identifying semantic boundaries, we can further compress the feature space of old objects as much as possible, thus reserving more network capacity for new objects, which can reduce feature conflicts. To meet this goal, we design a compact feature regularization, which helps us to eliminate the redundant features of old objects and leave more capacity for future unknown new objects.
Specifically, given a training sample, the latent features $M\in\mathbb{R}^{B\times C\times H\times W}$ is firstly compressed by aggregating spatial features as in Eq. (5),
@@
$S_{n}$ reflects the importance of semantic information for $M$ on each channel dimension. Larger eigenvalues correspond to more important semantic space, while smaller eigenvalues correspond to relatively non-primary semantic space [21, 13, 12]. Based on this fact, we can compress the features of each object by reducing the non-primary semantic information. Therefore, we construct the Semantic Compression Loss (SCL) as in Eq. (7) and Fig. 5,
@@
where $t\in[1,C]$ is a hyperparameter, which is a scale factor that reflects the degree of compression of the semantic feature space. To construct the total loss, we use $L=\lambda_{0}L_{1}+\lambda_{1}L_{CE}+\lambda_{2}L_{sc}$. $\lambda_{0}$ is set to 1 as a standard of the base task, and $\lambda_{1}$ is 0.5 as the auxiliary task experience for balancing different loss values. $\lambda_{2}$ is within the range of $1-10$, and in practice, users can make custom adjustment of $\lambda_{2}$ for controlling the old objects’ feature space. This loss continuously compacts the non-primary feature space while achieving the final optimization goal.
@@ 5 Experiments
5.1 Experimental Setup
Task Protocol According to our framework in Fig. 1 (E), our reconstruction model incrementally learns new objects. Based on the practical requirements of industrial defect inspection, we set up our experiments in both single-step and multi-step settings.
We represent our task stream as $\mathbf{X-Y\ with\ N\ Step(s)}$. Here, $\mathbf{X}$ denotes the number of base objects before starting incremental learning, $\mathbf{Y}$ represents the number of new objects incremented in each step, and $\mathbf{N}$ indicates the number of tasks during incremental learning. When training on base objects, $\mathbf{N}=0$, and after one step, $\mathbf{N}=\mathbf{N}+1$. Our task stream is shown as follows:
[…]
Table 1: Quantitative evaluation in MvTec AD [3] (A) and VisA [43] (B). “Image / Pixel” shows the image-level and pixel-level performance respectively. “NA” is Not Available, since CAD [20]-based method cannot locate defects in pixel-level. Red and Gray Red represents the best image-level and pixel-level performance respectively.
@@ 5.1 Experimental Setup
Figure 6: Qualitative evaluation in MVTec-AD [3] and VisA [43]. The intensity of red in the heatmap indicates a higher likelihood of defects, whereas blue signifies a lower probability. Our approach outperforms existing baselines by significantly mitigating semantic feature conflict and enhancing defect localization accuracy. (Please zoom in for more details.)
@@ 5.3 Ablation Study
Effectiveness of Object-Aware Self-Attention To demonstrate the effectiveness of category labels, we ablate Object-Aware Self-Attention in our structure. “w/o OASA” in Table 2 shows that this structure can reduce catastrophic forgetting by category information insertion. Moreover, in Fig. 7 (“w/o OASA”), if there is no semantic boundary, the network can still detect anomalies in the current object, but serious feature conflicts will occur in other regions of the object, further affecting the overall performance of the network.
Table 2: Quantitative evaluation in ablation study. The results show that three components, Object-Aware Self-Attention, semantic compression loss, and the updating strategy, contribute to the performance improvement of the whole framework. Red and Gray Red represents the best image-level and pixel-level performance respectively.
@@
Figure 7: Qualitative evaluation of ablation study. Compared with “w/o US” and “w/o OSOA”, it is obvious that our method can be more accurate in locating defects. Compared with “w/o SCL”, since increasing the available space of the model, our algorithm can further reduce the feature conflict and reduce more interference in the background. (Please zoom in for more details.)
Effectiveness of Semantic Compression Loss To demonstrate the effectiveness of Semantic Compression Loss, we ablate this loss in our training. “w/o SCL” in Table 2 shows that this structure can reduce catastrophic forgetting by preserving more space for new objects. Besides, in Fig. 7 (“w/o SCL”), by compressing more space for new objects, we can obtain clearer defect localization in detail, which indicates that more network space can be explained in incremental learning by semantic compression loss.
Effectiveness of Our Updating Strategy To demonstrate the effectiveness of our updating strategy, we ablate this updating method, and “w/o US” in Table 2 shows that this method can reduce catastrophic forgetting. Also, in Fig. 7 (“w/o US”), our algorithm can help memorize important semantic spaces for old objects, thus reducing the rewriting of semantic space in the reconstruction network, thus leading to a better localization performance.